### v1

# Upper bounds of focusing light through multimode fibers

Abstract
Wavefront shaping enables precise control of light propagation through multimode fibers, facilitating diffraction-limited focusing for applications such as high-resolution single-fiber imaging and high-power fiber amplifiers. While the theoretical intensity enhancement at the focal point is dictated by the number of input degrees of freedom, practical constraints—such as phase-only modulation and experimental noise—impose significant limitations. Despite its importance, the upper bounds of enhancement under these constraints remain largely unexplored. In this work, we establish a theoretical framework to predict the fundamental limits of intensity enhancement with phase-only modulation in the presence of noise-induced phase errors, and we experimentally demonstrate wavefront shaping that approaches these limits. Our experimental results confirm an enhancement factor of 5,000 in a large-core multimode fiber, approaching the theoretical upper bound, enabled by noise-tolerant wavefront shaping. These findings provide key insights into the limits of phase-only control in multimode fibers, with profound implications for single-fiber imaging, optical communication, high-power broad-area fiber amplification, and beyond.
pacs:
Valid PACS appear here
I Introduction
Multimode fibers (MMFs) are essential for high-resolution and ultra-thin single-fiber endoscopic imaging Čižmár and Dholakia (2012); Choi et al. (2012); Papadopoulos et al. (2012, 2013); Loterie et al. (2015); Amitonova and De Boer (2018); Amitonova and de Boer (2020), optical manipulation Čižmár and Dholakia (2011); Leite et al. (2018), high-bandwidth short-distance optical communication Li et al. (2021a), precise laser-based material processing Kakkava et al. (2019), and power scaling in fiber amplifiers Dawson et al. (2008); Richardson et al. (2010); Zervas and Codemard (2014); Fu et al. (2017). Their ability to support a large number of spatial modes increases the capacity for information transmission and energy delivery but also introduces challenges due to modal dispersion and complex interference effects. Overcoming these challenges requires precise control over light propagation within MMFs. Wavefront shaping provides a powerful approach to manipulate interference at the fiber’s output by tailoring the input field, enabling applications such as high-resolution imaging, targeted light delivery, and nonlinear effect management Di Leonardo and Bianchi (2011); Mosk et al. (2012); Horstmeyer et al. (2015); Vellekoop (2015); Rotter and Gigan (2017); Tzang et al. (2018); Gigan et al. (2022); Cao et al. (2022, 2023); Chen et al. (2023); Wisal et al. (2024a). Precise control of the output field through input wavefront shaping requires accurately measuring the fiber’s transmission matrix, which characterizes the input-output field relationship Popoff et al. (2010); Čižmár and Dholakia (2011); Choi et al. (2012); Kim et al. (2015); Li et al. (2021b); Bender et al. (2023). This measurement is essential for optimizing wavefront shaping techniques, including the fundamental task of focusing light at a desired location Vellekoop and Mosk (2007); Vellekoop et al. (2010); Papadopoulos et al. (2012); Caravaca-Aguirre et al. (2013); Amitonova et al. (2016); Descloux et al. (2016); Lyu et al. (2022); Gomes et al. (2022); Lyu and Amitonova (2024).
A key metric for evaluating wavefront shaping performance for focusing light through complex media is the enhancement factor, which quantifies the intensity at the focal position Vellekoop and Mosk (2007); Amitonova et al. (2016). The enhancement factor is defined as $\eta_{m}=I_{m}^{(\mathrm{foc})}/\langle I_{m}^{(\mathrm{rand})}\rangle$, where $I_{m}^{(\mathrm{foc})}$ is the focus intensity at the target position $m$ after wavefront shaping and $\langle I_{m}^{(\mathrm{rand})}\rangle$ is the averaged intensity at the same position $m$ over multiple random wavefront inputs Vellekoop and Mosk (2007); Vellekoop (2015); Mastiani et al. (2024). To maximize the enhancement factor, an optimization algorithm is typically used to determine the ideal input wavefront Vellekoop and Mosk (2008a); Yılmaz et al. (2013); Mastiani and Vellekoop (2021); Mastiani et al. (2024). Fundamentally, these algorithms are based on some form of transmission matrix measurement. In principle, with complete control over both the amplitude and phase of the input channels, all transmitted power can be efficiently collected at a desired focal position Gomes et al. (2022). However, in practice, there are certain limitations to the focus intensity due to incomplete input channel control Resisi et al. (2020); Mastiani et al. (2024). With phase and amplitude modulation of the input wavefront, the enhancement factor is equal to the number of input degrees of freedom denoted as $\eta=N$. Nevertheless, most spatial light modulators (SLM) function primarily as phase-only modulators, as amplitude modulation is generally avoided in practice Mosk et al. (2012). This is because amplitude modulation reduces the input power, making it less efficient for applications requiring high-intensity light focusing, such as broad-area fiber amplifiers and laser ablation through large-core optical fibers Kakkava et al. (2019); Wisal et al. (2024b); Rothe et al. (2025a, b). Therefore, in practical scenarios, phase-only input modulation is typically preferred; the input wavefronts can be optimally shaped to achieve constructive interference at the desired focus location at the fiber’s output, maintaining a constant input power Gigan et al. (2022); Cao et al. (2023); Vellekoop and Mosk (2008a); Rothe et al. (2025a). This constraint inherently limits the maximum enhancement factor, with its theoretical upper bound, following $\eta=R(N-1)+1$, where $R=\pi/4$ represents the participation ratio—a well-established result for focusing light through disordered scattering media. Vellekoop and Mosk (2007); van Beijnum (2009); Vellekoop and Mosk (2008a); Yılmaz et al. (2013); Vellekoop (2015); Mastiani et al. (2024).
Although the assumption that the participation ratio is $R=\pi/4$ has been widely used to estimate the upper bound of the enhancement factor in wavefront shaping through multimode fibers (MMFs), its validity has not been critically examined. This raises a fundamental question: is the participation ratio for phase-only modulation in MMFs $R=\pi/4$ similar to that in disordered media? If so, what is the upper limit of the achievable enhancement factor, and can it be experimentally reached?
Here, we present a comprehensive work that combines both experimental and theoretical aspects of wavefront shaping. We introduce the theoretical upper bound on the enhancement factor for phase-only modulation; moreover, we experimentally demonstrate it for focusing light through a multimode fiber (MMF). We first describe how the participation ratio $R$ depends on the basis used for phase-only modulation. Our theoretical calculations and experimental observations reveal that when wavefront modulation is performed on the Fourier space of the proximal end of the fiber, the $R_{m}=\pi/4$ at any fiber output position $m$, which closely resembles the well-known result observed in disordered media. In contrast, phase-only modulation in the fiber mode basis reveals a strong radial dependence of $R$ at the fiber’s distal end. To experimentally approach the theoretical upper bound of the enhancement factor for phase-only modulation, we perform noise-tolerant transmission matrix measurements using a Hadamard basis, achieving an enhancement factor of $\eta=5,000$. We also introduce a predictive method that combines our theoretical framework with measured phase errors in a practical wavefront shaping setup. Using this approach, we quantify the phase errors in transmission matrix measurements both in the canonical (SLM pixel) and Hadamard bases. This method provides a clear quantitative understanding of how Hadamard-based measurements minimize phase errors, enabling near-ideal wavefront shaping. By linking theoretical predictions with experimental observations, this work establishes a foundation for accurately predicting the enhancement factor and achieving near-perfect phase-only wavefront shaping in multimode fibers. Our findings directly advance the development of more robust and efficient wavefront shaping techniques, with broad applications in high-resolution imaging, broad-area fiber amplifiers, laser ablation through multimode fibers, and beyond.
Figure 1: Wavefront shaping setup and the results are shown. (a) Experimental setup: The spatial light modulator (SLM) modulates the laser beam on the multimode fiber’s proximal end and focuses on the distal end. P: linear polarizer; HWP: half-wave plate; BS: beam splitter; M: mirror; ${\rm MO}_{1}$ and ${\rm MO}_{2}$: microscope objectives; MMF: multimode fiber; NA: numerical aperture; CCD: charge-coupled device. (b) An experimental image of the speckle formation at the distal end of the MMF is shown when a random wavefront is incident on the proximal end. The interference of the waves propagating through various optical modes in the MMF results in random intensity fluctuations, giving rise to the granular appearance of speckle patterns. Here, the fiber radius is $a=100$ µm. (c) Experimental image of the distal end of the fiber when light is focused by wavefront shaping on the canonical (SLM pixel) basis and (d) on the Hadamard basis with $N=8,000$. The scale bar indicates the intensity across the distal end as observed on the CCD camera and is normalized to the highest count on the image. (e) The mean enhancement factor $\eta$ averaged over azimuthal $\theta$ positions versus the normalized radial distance $r/a$ at the fiber distal end for the number of degrees of freedom $N=8,000$. The blue solid line represents the upper limits of the enhancement factor with full-field modulation at the input equal to $\eta=N$. The violet solid line represents the upper bounds of the enhancement factor with perfect phase-only modulation when the SLM is placed on the Fourier plane of the fiber proximal end. The black and red solid lines represent the experimental enhancement factors with wavefront shaping on the Hadamard and canonical (SLM pixel) basis. The enhancement factor is higher with wavefront shaping on the Hadamard basis.
II Transmission matrix and focusing measurements
The first step in controlling light propagation through a multimode fiber (MMF) involves measuring its transmission matrix. The transmission of light through an MMF can be described with a transmission matrix with elements $t_{mn}$
$E_{m}=\sum_{n=1}^{N}t_{mn}E_{n}$ (1)
where $m$ and $n$ are the indices of the outgoing and incident fields, and $N$ is the number of independently controlled degrees of freedom at the input.
Using the spatial light modulator in our experimental setup, shown in Fig. 1(a), we select an input basis indexed by $n$ (either canonical or Hadamard) for the incident fields $E_{n}$, as defined in Equation 1. The field transmission matrix is then measured using common-path phase-shifting interferometry Popoff et al. (2010); Hsu et al. (2017); Yılmaz et al. (2019, 2019) in a chosen basis.
Figure 2: (a) The mean participation ratio $R$, averaged over the core radius $a$ and the number of degrees of freedom $N$, versus the normalized radial distance $r/a$. The experimental $R$ (solid red line) is consistent with the numerical $R$ (solid blue line) in the Fourier basis, both following a trend close to $\pi/4$ (dashed black line) and showing no dependence on the radial distance. However, we observe a strong dependence of $R$ on the radial distance when the $R$ is computed with phase-only modulation on the MMF fiber mode basis (solid black line). (b) The mean participation ratio $R$, averaged over the core radius $a$ and the radial distance $r$ at the fiber distal end, shown with respect to the normalized number of degrees of freedom $N/N_{\text{mode}}$. The experimental $R$ (red line) agrees with the numerical $R$ (solid blue line) and remains invariant with $N/N_{\text{mode}}$, maintaining a value near $\pi/4$ (dashed black line). (c) The mean enhancement factor $\eta$, averaged over the azimuthal position $\theta$ and radial distances $r$ at the fiber distal end, shown with respect to the number of degrees of freedom $N$ = 172, 484, 952, and 2,032. The experimental enhancement factor $\eta$ in the Hadamard basis (red solid line) closely follows the theoretical prediction (black solid line) and is notably higher than the experimental $\eta$ in the canonical basis (blue solid line).
To measure the transmission matrix of the MMF using common-path interferometry Popoff et al. (2010), the light field on the spatial light modulator is divided into a signal and a reference part. The signal is modulated with four phase steps uniformly spaced between 00 and $2\pi$, and both the signal and reference components are then coupled into the MMF. The MMF is a step-index fiber with a diameter of 200 µm, a length of 6 cm, and an NA = 0.22. The light propagates through the supported modes of the MMF, and the resulting intensity pattern at the distal end is imaged onto a charged-coupled device (CCD) camera sensor. In this work, we use feedback-based wavefront shaping to retrieve the transmission matrix from the SLM (in the Fourier space) to the CCD camera (in the real space) using two different algorithms: the stepwise sequential algorithm (SSA) on the canonical basis Vellekoop and Mosk (2007) and the dual reference algorithm on the Hadamard basis Mastiani and Vellekoop (2021). For both methods, we divide the SLM surface into $N$ number of input degrees of freedom (segments) where $N$ is chosen to be 172, 484, 952, 2,032, 3,300, 3,940, 5,388, 6,180, 7,080, and 8,000 in all the experiments described here. We choose $N$ in such a way that we display a circular phase pattern on the SLM. In our experimental setup, depicted in Fig. 1(a), we modulate a single linear horizontal polarization of light at the input and detect the same linear horizontal polarization at the output (the details are described in the supplementary material, section I - A).
Maximizing the overlap between the input field patterns and the fiber’s core at its proximal end, both in real and Fourier space, is essential. In real space, the input field patterns must be centered at the fiber core, and their size should be equal to or smaller than the fiber core diameter. In Fourier space, the diameter of the displayed SLM pattern should correspond to a numerical aperture equal to or smaller than the fiber’s numerical aperture, NA = 0.22. We adjusted our setup such that displaying SLM patterns with $6\times 6$ pixel-size segments ensures a near-perfect size match between any arbitrary SLM pattern and the fiber core (additional details can be found in the supplementary material, section I - A). Furthermore, the diameter of the SLM pattern consistently corresponds to an NA smaller than that of the fiber, with NA $<$ 0.22.
The phases of each transmission matrix column $m$ are computed using four recorded CCD camera images, each corresponding to a different relative phase between the signal and the reference, obtained through common-path four-phase shifting interferometry with the SLM Popoff et al. (2010); Hsu et al. (2017); Yılmaz et al. (2019, 2019). The transmission matrix elements map the input field onto the output field, where every element on the output field is a summation of all the input elements multiplied by an analogous transmission coefficient. Thus, a random summation of $N$ field components contributes to each element in the output field, which results in the speckle pattern as is seen in Fig. 1(b). To form a focus, i.e., to increase light intensity for a specific output element, the $N$ field components must constructively interfere with each other. This is accomplished by displaying the conjugate of the measured transmission matrix phase row with index $m$, which corresponds to the desired position of the output field, on the SLM.
[…]
In the dual reference algorithm Mastiani and Vellekoop (2021), we use the Hadamard basis, where the size of the basis must be $N_{1}=2^{p}$, with $p$ being an integer (e.g., $N_{1}$ = 128, 256, 512, 1,024, 2,048, 4,096). In this approach, we divide the SLM segments into two equal-sized groups, each containing a small number of overlapping segments, denoted as $O$. In the first step of the algorithm, a Hadamard pattern is displayed on the segments of group 1 (segments from 1 to $N_{1}$), while the remaining segments (from $N_{1}+1$ to $N$) contribute as the reference field $E_{\rm ref}^{(1)}$. Four-phase shift interferometry is applied to both groups to vary their relative phase from 0 to $2\pi$ Popoff et al. (2010). Subsequently, we perform a Hadamard transform on the full-field patterns obtained from the four-phase-shifting interferometry to extract the transmission matrix phase elements for group 1 relative to the $E_{\rm ref}^{(1)}$ phase, which results in the transmission matrix 1 $t^{(1)}$.
In the second step of the algorithm, the segments of group 2 (from $N-N_{1}+1$ to $N$) are modulated, while the remaining segments (from 1 to $N-N_{1}$) serve as reference $E_{\rm ref}^{(2)}$. The procedure is then repeated, producing two matrices, $t^{(1)}$ and $t^{(2)}$, with phase values relative to the two different reference fields $E_{\rm ref}^{(1)}$ and $E_{\rm ref}^{(2)}$ for each camera pixel $m$. To obtain the final transmission matrix, we need to determine the phase difference between $E_{\rm ref}^{(1)}$ and $E_{\rm ref}^{(2)}$ at each camera pixel $m$. By using the phase differences between the overlapping segments $O$ relative to $E_{\rm ref}^{(1)}$ and $E_{\rm ref}^{(2)}$ in both steps of the algorithm, we calculate the phase difference between $E_{\rm ref}^{(1)}$ and $E_{\rm ref}^{(2)}$. Finally, we adjust the phases of $t^{(1)}$ and $t^{(2)}$ based on this known phase difference and combine the two matrices to construct the final full-field transmission matrix.
Note that, in all transmission matrix experiments, we recorded fiber output intensity images $I_{m}^{(\mathrm{rand})}$ for 1,000 random wavefront inputs, using the same number of segments $N$ as in each experiment. Fig. 1(b) shows an example random speckle pattern observed on the CCD camera for a random input wavefront. These speckles result from modal dispersion, where each fiber mode propagates with a distinct propagation constant and phase delay, leading to complex interference at the output.
Immediately after each transmission matrix measurement, we perform focusing experiments. To focus at a specific position $m$ within the core at the fiber’s distal end, we display the conjugated phase from the measured transmission matrix row corresponding to position $m$ and record the focused intensity pattern on the CCD camera. To avoid CCD camera saturation, calibrated neutral density (ND) filters were placed in front of the camera during the measurements. These experiments are performed to focus on various radial $(r)$ and azimuthal $(\theta)$ positions within the core of the fiber’s distal end using phase-only input modulation, obtained from the corresponding transmission matrix data. The resulting enhancement factors are calculated using the expression $\eta_{m}=I_{m}^{(\mathrm{foc})}/\langle I_{m}^{(\mathrm{rand})}\rangle$, where $I_{m}^{(\mathrm{foc})}$ represents the focused intensity at the target position $m$ after wavefront shaping, and $\langle I_{m}^{(\mathrm{rand})}\rangle$ represents the averaged intensity at the same position $m$ over 1,000 random wavefront inputs. Two example intensity patterns are shown in Fig. 1(c) and Fig. 1(d) for wavefront shaping on the canonical and Hadamard basis, respectively. As evident from the captured images, the optimized focus achieved using the Hadamard basis exhibits a higher peak intensity compared to the focus obtained on the canonical basis.
Theoretically, for full-field modulation, the expected enhancement factor is given by $\eta=N$, where $N$ represents the number of input degrees of freedom. Thus, for $N=8,000$, the theoretical full-field enhancement is $\eta=8,000$. Our numerical simulations confirm that the expected prefactor for phase-only modulation is $\pi/4$, leading to a theoretical enhancement factor of $\eta=6,135$ for $N=8,000$.
Fig. 1(e) presents the enhancement factor $\eta$ as a function of the normalized radial distance $r/a$, comparing wavefront shaping using the transmission matrix measured in the canonical and Hadamard bases. The results clearly show that using the Hadamard basis yields a higher enhancement factor across all radial distances within the core at the fiber’s distal end.
Figure 3: Phase error measurements in wavefront shaping experiments. (a) A conceptual sketch of two independently measured phase maps, $\phi_{mn}^{(1)}$ and $\phi_{mn}^{(2)}$ for the same input $n$. The phase difference, $\delta\phi_{mn}^{\prime}=\arg\left(e^{i(\phi_{mn}^{(2)}-\phi_{mn}^{(1)})}\right)$, represents the measured phase error. (b) Standard deviation $\sigma\mathrm{{}_{\delta\phi}}$ of the Gaussian-fitted phase error distributions as a function of the number of degrees of freedom $N$. In the canonical basis, $\sigma\mathrm{{}_{\delta\phi}}$ increases with $N$, indicating higher phase errors. In contrast, the Hadamard basis maintains a consistently low $\sigma\mathrm{{}_{\delta\phi}}$, suggesting greater robustness to phase errors. (c, d) Probability density functions of the phase errors $P(\delta\phi_{mn}^{\prime})$ for the canonical and Hadamard bases, respectively, as a function of $N$. The Gaussian-fitted curves (red) show that in the canonical basis, the phase error distribution broadens significantly with increasing $N$. In contrast, (d) shows that the Hadamard basis maintains a sharply peaked distribution around zero, indicating minimal phase errors.
Figure 4: The impact of phase errors on the enhancement factor. (a) Schematic representation of the phase error coefficient $\Phi$ for the Hadamard (red) and canonical (blue) bases, where $\delta\phi$ represents phase errors. The dashed lines indicate how phase errors accumulate differently in the two bases, with the Hadamard basis maintaining $\Phi\approx 1$ due to a reduced cumulative effect of phase errors. (b) The phase error coefficient $\Phi$ as a function of the number of degrees of freedom $N$. On the canonical basis, $\Phi$ decreases with $N$ due to a decreasing signal-to-noise ratio (SNR), which reduces interferometric visibility. In contrast, the Hadamard basis maintains $\Phi\approx 1$ due to a balanced signal-to-background ratio. (c) The phase error coefficient $\Phi$ for $N=8000$, plotted as a function of the normalized photon budget. The Hadamard basis maintains a consistently high $\Phi$ compared to the canonical basis, reflecting its improved SNR under the same noise conditions. The normalized photon budget (ranging from 0 to 1) is defined so that a value of 1 corresponds to a maximum photon budget of approximately 214 mean counts on the CCD.
III The enhancement factor and its practical limitations
We present a general form of the enhancement factor at the output $m$ in terms of two major parameters; the input participation ratio $R$ and the phase error coefficient $\Phi$:
$\eta_{m}=\alpha R_{m}\Phi_{m}(N-1)+1,$ (2)
where $\alpha\equiv\left\langle A_{n}\right\rangle_{n}^{2}/\left\langle A_{n}^{2}\right\rangle_{n}$ defines the spatial homogeneity of the amplitude profile incident on the SLM, $R_{m}\equiv\left\langle|t_{mn}|\right\rangle_{n}^{2}/\left\langle|t_{mn}|^{2}\right\rangle_{n}$ quantifies the input participation ratio at the output $m$, $\Phi_{m}\equiv\left\langle{\rm cos}(\delta\phi_{mn})\right\rangle_{n}^{2}$ quantifies the phase error coefficient at the output $m$, and $N$ represents the number of controlled degrees of freedom at the input wavefront. Here $\left\langle\right\rangle_{n}$ denotes averaging over inputs $n$. The phase error coefficient $\Phi$ varies between 0 and 1, where, in the absence of phase errors, it equals $\Phi=1$ van Putten (2011). In our experiments, $\alpha=1$, since we illuminate the SLM with a flat top laser beam expanded from a single-mode fiber output. In our experimental system, the input index $n$ corresponds to the SLM segment index (approximately the Fourier components at the proximal end) and the output index $m$ to the CCD camera pixels (position at the distal end), respectively. Note that the enhancement factor is defined differently in certain studies Gomes et al. (2022); Duan et al. (2023), whereas the enhancement factor we establish here is standardized Vellekoop and Mosk (2007, 2008b, 2008a); Papadopoulos et al. (2012); Yılmaz et al. (2013); Vellekoop (2015); Amitonova et al. (2016); Resisi et al. (2020); Mastiani et al. (2024) and possesses a clearly defined upper limit $\eta=N_{\rm mode}$ with complete ($N=N_{\rm mode}$) full-field (amplitude and phase) input modulation.
III.1 The enhancement factor for phase-only input modulation
In phase-only wavefront shaping experiments, evaluating the input participation ratio $R_{m}$ is essential, as it quantifies the fraction of input degrees of freedom that effectively contribute to the output at position $m$ under phase-only modulation. We quantify $R_{m}$ both in experiments and numerical simulations. To obtain the numerical participation ratio, we first compute the transmission matrix using the mode decomposition method Lee (2023) (see Section II in the supplementary material for details). To evaluate the participation ratio in the desired input basis, we apply a basis transformation to the columns of the transmission matrix and compute the participation ratio in the transformed basis. To obtain the experimental input participation ratio, we complemented phase measurements of the transmission matrix—acquired using the dual reference algorithm in the Hadamard basis—with amplitude measurements by coupling only the signal component of the light from the SLM into the fiber. This approach allowed us to reconstruct both the amplitude and phase of the transmission matrix elements $t_{mn^{\prime}}$, where $n^{\prime}$ represents the Hadamard vector index. We then applied a Hadamard transform as $t_{mn}=\sum_{n^{\prime}=1}^{N}t_{mn^{\prime}}H_{n^{\prime}n}$ to convert the transmission matrix into the canonical (SLM pixel) basis $n$. Here $H_{n^{\prime}n}$ represents the unitary Hadamard transform matrix. The participation ratio was subsequently calculated on this basis, as our phase-only modulation experiments are performed in canonical bases.
Fig. 2(a) presents both numerical and experimental participation ratios $R$ as a function of the normalized radial distance $r/a$. The numerical results in the Fourier basis closely align with the experimental data, exhibiting the same overall trend. Notably, when phase-only input modulation is performed on the Fourier basis, $R$ remains constant regardless of the focal position at the distal end of the fiber. Additionally, placing the spatial light modulator (SLM) in the Fourier plane of the MMF’s proximal end yields a participation ratio of $R=\pi/4$, consistent with the well-known value observed in phase-only wavefront shaping through random media van Beijnum (2009). In the experiments, the number of guided modes is $N_{\rm mode}=15,178$ for a fiber with a core radius of $a=100$ µm and numerical aperture (NA) of 0.22. In the simulations, we consider four fibers with core radii of $a$ = 15 µm, 22 µm, 28 µm, and 38 µm (all with NA = 0.22), corresponding to $N_{\rm mode}$ = 180, 374, 606, and 1,114, respectively. All numerical values are calculated for light with a wavelength of 561 nm.
While implementing phase-only modulation directly on the fiber mode basis is not practical, it offers valuable insights into the fundamental behavior of phase-only wavefront shaping. For this reason, we also numerically calculated the participation ratio on the fiber mode basis. Our simulations reveal that when phase-only modulation is applied on this basis, the participation ratio $R$ shows a strong dependence on the radial position of the focal point at the distal end of the MMF. At the center of the fiber core, the participation ratio drops below 0.2, as most fiber mode wavefunctions exhibit ring-shaped profiles and contribute minimally at the core center.
Fig. 2(b) shows the radially averaged theoretical and experimental participation ratios $R$ (both for Fourier basis) as a function of the normalized number of degrees of freedom $N/N_{\text{mode}}$, where the experimental $N_{\text{mode}}$ is determined from the equation $V=2\pi a\text{NA}/\lambda$, with $\lambda=561$ nm, $a=100$ µm (core radius), and NA = 0.22. For our fiber, the number of modes per polarization is $N_{\text{mode}}=V^{2}/4=15,178$. Our experimental and numerical results consistently show that when phase-only input modulation is performed on the Fourier basis, the participation ratio remains $R=\pi/4$, regardless of the number of input degrees of freedom.
Fig. 2(c) presents the enhancement factor $\eta$, averaged over radial ($r$) and azimuthal ($\theta$) coordinates, as a function of $N$. We observe that the experimental enhancement factor closely approaches the theoretical limit without phase errors, $\eta=(\pi/4)(N-1)+1$, when the transmission matrix is measured on the Hadamard basis. For $N=2,000$, the enhancement factor obtained in the canonical basis is lower than the theoretical prediction, as the signal-to-noise ratio decreases with increasing $N$, leading to phase errors ($\Phi<1$).
Our theoretical and experimental results demonstrate that the widely recognized $\pi/4$ factor observed in wavefront shaping through disordered media also holds for wavefront shaping through multimode fibers. This similarity arises because each degree of freedom in the Fourier space couples randomly to multiple superpositions of fiber modes, leading to output speckle statistics that closely resemble those seen in disordered media.
III.2 The effect of phase errors on the enhancement factor
Accurate transmission matrix measurements are essential to approach the theoretical upper bound of the enhancement factor. However, experimental noise introduces phase errors that degrade these measurements Vellekoop and Mosk (2008a); Yılmaz et al. (2019). Here we introduce a predictive methodology to quantify these phase errors by applying the phase-shift interferometry measurement twice for the same SLM input pattern. In the absence of phase errors, the measured phases $\phi_{mn}^{(1)}$ and $\phi_{mn}^{(2)}$ must be identical for the same input $n$ and output $m$. However, in practice, experimental noise gives rise to phase errors, $\delta\phi_{mn}^{(1)}$ and $\delta\phi_{mn}^{(2)}$, in both measurements. We extract these phase errors as $\delta\phi_{mn}^{\prime}=\arg\left(e^{i(\phi_{mn}^{(2)}-\phi_{mn}^{(1)})}\right)$. A sketch is shown in Fig. 3(a) to describe the phase error measurement concept. Assuming the phase errors follow a Gaussian distribution with zero mean—which is a reasonable assumption as evidenced by Figs. 3(c) and Figs. 3(d)—we estimate the phase errors $\delta\phi_{mn}$ in a single four phase shift interferometry measurement (see section I-D in the supplementary material for details).
Fig. 3(b) presents the standard deviation $\sigma_{\delta\phi^{\prime}}$ of the least-squares fitted Gaussian functions to the experimental phase error histograms for different numbers of degrees of freedom $N$. The Hadamard basis maintains a consistently low $\sigma_{\delta\phi^{\prime}}$, indicating that phase errors remain minimal variations. In contrast, the canonical basis exhibits an increasing $\sigma_{\delta\phi^{\prime}}$ as $N$ grows, revealing a broadening phase error distribution. This trend suggests that phase errors in the canonical basis become more pronounced with an increasing number of degrees of freedom, leading to greater deviations from the ideal constructive interference.
Wavefront shaping relies on precise constructive interference at the target position $m$, which makes it sensitive to experimental noise that introduces phase errors. To characterize the level of constructive interference in our experiments, we introduce the phase error coefficient $\Phi_{m}\equiv\left\langle{\rm cos}(\delta\phi_{mn})\right\rangle_{n}^{2}$ at the target position $m$. This coefficient equals unity in the absence of phase errors and approaches zero when the phase measurements contain no usable information. Fig. 4(a) illustrates how phase deviations $\delta\phi_{mn}$ disrupt the alignment of field contributions, reducing $\Phi$ at the output position $m$. In an ideal wavefront shaping experiment, the phases of the individual field contributions are perfectly aligned, leading to strong constructive interference and high intensity at the target. In practical, high-quality experiments, small phase errors reduce this alignment slightly, resulting in $\Phi$ values close to 1. However, in canonical basis measurements, $\Phi$ is generally much lower—especially for large $N$—due to more significant phase mismatches.
To quantify the impact of accumulated phase errors, we experimentally quantified the phase error coefficient $\Phi$ at each output position $m$ and computed its ensemble-averaged value $\langle\Phi_{mn}\rangle_{n}$ as a function of the number of degrees of freedom $N$, as shown in Fig. 4(b). In the Hadamard basis, $\Phi$ remains close to 1, while in the canonical basis, $\Phi$ decreases significantly as $N$ increases. This degradation is due to the fact that Hadamard wavefront shaping maintains a high signal-to-noise ratio by balancing the signal and reference fields, ensuring maximum interferometric visibility. In contrast, on the canonical basis, as $N$ increases, the signal diminishes, reducing the signal-to-noise ratio, degrading interference visibility, and limiting the accuracy of the measured transmission matrix.
To further explore how the photon budget influences phase errors, we examine the behavior of $\Phi$ for a fixed $N=8000$ as a function of the incident power on the CCD camera (seen in Fig. 4(c)). The incident power is adjusted by placing neutral density (ND) filters in front of the CCD chip. The Hadamard basis maintains $\Phi\approx 1$ even as the available photon budget decreases, demonstrating its robustness to photon-limited conditions Jang et al. (2017). This stability arises because Hadamard-based measurements preserve high interferometric contrast and balanced signal-to-reference ratios, maintaining a reliable signal-to-noise ratio for phase retrieval. However, on the canonical basis, $\Phi$ steadily drops as the photon budget decreases. The degradation stems from a reduced signal-to-noise ratio, leading to increased phase errors. As expected, when the signal photon count approaches zero, phase error coefficient estimation deteriorates for both bases, ultimately preventing phase retrieval.
IV Results and Discussion
Figure 5: The mean enhancement factor $\eta$, averaged over the azimuthal $\theta$ positions and radial distance $r$ at the fiber’s distal end, as a function of the number of degrees of freedom $N$. Solid blue and red lines show the theoretical enhancement factors, while dashed lines represent experimental measurements for wavefront shaping on the canonical and Hadamard bases, respectively. Error bars indicate the standard deviation of enhancement factors measured at different focal positions on the fiber’s distal end.
Here, we report the enhancement factor values for focusing at various radial distances ($r$) and azimuthal angles ($\theta$) within the fiber core at the distal end, for different values of $N$, using transmission matrices measured in both the canonical and Hadamard bases. Additionally, we provide the predicted enhancement factors obtained from our method, which uses the measured phase error coefficients $\Phi$ for both bases.
In Fig. 5, we display the measured and predicted enhancement factor $\eta$ as a function of the number of degrees of freedom, $N$. Transmission matrix measurements in the Hadamard basis result in improved enhancement factors, thanks to its superior signal-to-noise ratio, particularly as $N$ increases. However, for $N<952$, the results from both the canonical and Hadamard bases are nearly identical, indicating that the advantages of the Hadamard basis become more pronounced as the number of degrees of freedom grows.
For lower values of $N$ ($N<3000$), the measured enhancement factor closely aligns with the expected theoretical values. For larger $N$, however, the measured enhancement deviates from the theoretical predictions, consistently remaining lower due to experimental limitations. Below $N=3{,}000$, phase errors caused by experimental noise are the primary limiting factor. Beyond this threshold ($N>3{,}000$), transmission matrix decorrelation becomes the dominant constraint, especially during measurement periods extending up to four hours.
V Conclusion
In conclusion, we present a methodology to predict the upper bound of the enhancement factor for focusing light through a multimode fiber using phase-only modulation. Our approach combines theoretical analysis with experimental phase error measurements obtained during practical transmission matrix acquisition. Using phase-only modulation of the incident wavefront, we experimentally approach this theoretical upper limit. Our results demonstrate, both theoretically and experimentally, that the participation ratio for phase-only input modulation in the Fourier domain at the fiber’s proximal end is $R=\pi/4$, consistent with the well-known result for disordered random media.
Furthermore, our method provides a quantitative explanation for the differences in enhancement factors observed in wavefront shaping experiments using Hadamard and canonical bases, attributing them to phase errors. This insight opens up new possibilities for optimizing wavefront shaping techniques. The upper limit we establish is particularly important for applications relying on phase-only modulation, such as laser ablation through large-core optical fibers Kakkava et al. (2019), nonlinear effect suppression and clean beam formation for broad-area fiber amplifiers Chen et al. (2023); Wisal et al. (2024a, b, c); Rothe et al. (2025a).
Acknowledgements.
We thank Mesut Laçin and F. Ömer İlday for their technical support for fiber preparation and A. Serhan Başdemirci for useful discussions. This work is financially supported by TÜBİTAK grant no. 122F311 and by the BAGEP Award of the Science Academy with funding supplied by Sevinç - Erdal İnönü Vakfı.